OpenStudy (anonymous):

Which equation is y = 2x2 – 8x + 9 rewritten in vertex form? y = 2(x – 2)2 + 9 y = 2(x – 2)2 + 5 y = 2(x – 2)2 + 1 y = 2(x – 2)2 + 17

OpenStudy (anonymous):

y = 2(x – 2)2 + 1

OpenStudy (anonymous):

\(\Large y = 2x^2 – 8x + 9=2(x^2-4x+\frac{9}{2})\\\Large =2[(x^2-2\times2x+2^2)+(\frac{9}{2}-2^2)] =2(x-2)^2+1\)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

:)